OpenStudy (anonymous):

Rank axis from smallest to largest: 1 4 -4

OpenStudy (anonymous):

axis of symmetry

OpenStudy (anonymous):

my order: 4, 1 , -4 is that right?

OpenStudy (mathmale):

Try plotting 1, 4 and -4 on the x-axis. Which comes first? second? third?

OpenStudy (mathmale):

You want them

OpenStudy (mathmale):

in increasing order of size.

OpenStudy (anonymous):

"rank the functions based on their axis of symmetry (from smallest to largest)" this is what the question asks @mathmale

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

@Serenity74

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ValentinesDayAngel

OpenStudy (anonymous):

@Directrix any thoughts?

Directrix (directrix):

Sorry. I've not heard of ranking functions by axis of symmetry. I need to study.

OpenStudy (mathmale):

Go back and check the wording of the original question. I agree with Directrix (above).

Directrix (directrix):

>>Try plotting 1, 4 and -4 on the x-axis. If you follow mathmale's hint, then your answer of >> my order: 4, 1 , -4 is not correct because you have the numbers in decreasing order of size. Reading the x-axis from left to right, 4 does not come before 1 and 1 does not come before -4. @Everly

Directrix (directrix):

Yes. Please post the functions associated with the axes of symmetry.

OpenStudy (anonymous):

Oh right: so -4 , 1 , 4 @Directrix ?

Directrix (directrix):

If the original but unposted functions are parabolas that open upward or downward, then the lines of symmetry x = -4, x = 1, and x = 4 would appear as parallel lines in that order. If I understand the question, then, yes, I think your answer is now correct. I can't say that I am certain, though. @Everly

OpenStudy (anonymous):

well this is the question

OpenStudy (anonymous):

These are my answers The axis of symmetry is simply the x coordinate of the vertex. g(x), x = -b/(2a) = 16/(2*2) = 16/4 x = 4, the axis of symmetry is 4. h(x), the vertex is (1, -3). Therefore the axis of symmetry is 1. f(x), the expression is in vertex form: a(x-h)^2 +k. The variables h and k represent the vertex so (-4, 1) The axis of symmetry is -4.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

The problem says to rank the functions based on each one's axis of symmetry from small to large. Rather than answer with the equations of the axes of symmetry, I think you are to respond with the functions. f(x), h(x), and g(x) That order corresponds to the -4, 1, 4 which you mentioned earlier in the post. @Everly

OpenStudy (anonymous):

Oh alright that sounds just about right, thank you so much !

Directrix (directrix):

You are welcome.